OpenStudy (anonymous):

Caroline runs 3 1/2 miles in 22 1/4 minutes. How many miles does she run per minute? A) 7/12 mile per minute B) 1 5/7 mile per minute C) 1 1/15 mile per minute D) 1 4/11 mile per minute

OpenStudy (anonymous):

=w= im not gonna know................... Grr.................... Looks like imi gettin' an f~

OpenStudy (anonymous):

(3 1/2) /( 22 1/4) = x/1 cross multiply (89/4)(x) = (7/2)(1) 89/4x = 7/2 x = (7/2)/ (89/4) x = 7/2 * 4/89 x = 28/178 = 14/89 I am not getting any of your answer choices Is the problem written correctly ?

OpenStudy (anonymous):

._. yes...

OpenStudy (anonymous):

I am sorry....I do not know

OpenStudy (anonymous):

its cool :/